OpenStudy (anonymous):

Find the solution to the differential equation dy/dx=√(xy^3) that satisfies the initial condition y(0)=4

OpenStudy (irishboy123):

separate it!

OpenStudy (anonymous):

@IrishBoy123 i got stuck after this

OpenStudy (irishboy123):

line 5 looks dodge! should get something like \(\sqrt{y} = \dfrac{1}{c - \sqrt{x^3}}\)as you work through it